# Towards Backdoor Stealthiness in Model Parameter Space

### v2

@@ 3.2. Lack of Parameter-Space Stealthiness
As introduced in the related work, early backdoor attacks that introduce noticeable changes in either input (Gu et al., 2019; Chen et al., 2017) or feature space (Nguyen and Tran, 2021, 2020) have been empirically shown powerful, even with very low poisoning rates (Gu et al., 2019; Zeng et al., 2023). Focusing on the backdoor-introduced noticeable changes, backdoor defenses are improved to distinguish backdoor patterns in either input or feature space (Wang et al., 2022a; Lin et al., 2024). Meanwhile, backdoor attacks are optimized to increase stealthiness in input (Nguyen and Tran, 2021) or feature space (Qi et al., 2023a). However, regardless of the implementation of input- or feature-space attack logic, backdoor behaviors are eventually embedded in the backdoored model’s parameters. For this reason, it is important to investigate whether backdoor attacks introduce visible changes in the parameter space of the attacked models that can be used by the parameter-space defenses.
Considering this observation, we ran an initial experiment to understand the behavior of neurons in backdoored models. We use the TAC values (Zheng et al., 2022) to quantify the relevance of a neuron to the backdoor behavior. The TAC values show the change in the output of each neuron (a feature channel of convolutional layers) before and after the trigger is attached to the input. Thus, TAC quantifies a neuron’s sensitivity to the backdoor trigger. A high TAC value indicates that the neuron is strongly related to the backdoor behavior, whereas a low TAC value shows it is not. TAC takes the exact trigger information into account when measuring the backdoor effect, which makes TAC a straightforward and effective method that captures the relevant backdoor neurons’ behaviour. Specifically, TAC is defined as:
(2) $\text{TAC}_{l}^{(k)}(\mathcal{D}_{c})=\frac{1}{|\mathcal{D}_{c}|}\sum_{{\bm{x}}\in\mathcal{D}_{c}}||f_{l}^{(k)}({\bm{x}})-f_{l}^{(k)}(G_{x}({\bm{x}}))||_{2},$
@@ 3.3. Grond for Comprehensive Stealthiness
Backdoor trigger generation for input-space stealthiness. We use imperceptible adversarial perturbations to generate imperceptible backdoor triggers inspired by adversarial example studies (Moosavi-Dezfooli et al., 2017; Zhang et al., 2021). We modify the original PGD algorithm to generate a universal PGD (UPGD) perturbation as the backdoor trigger. UPGD contains non-robust but generalizable semantic information (Tsipras et al., 2019), which correlates with the benign functions of the victim model and shortens the distance between poisoned data and the target classification region (Zhang et al., 2021). Consequently, backdoor patterns tend to make fewer prominent changes to the victim network.
Similar to (Zeng et al., 2023; Zhang et al., 2021), UPGD is generated on a well-trained surrogate model trained on the clean training set. The architecture and parameters of the surrogate model do not necessarily need to be the same as the victim model (see Table 15 in App. C.2). UPGD is optimized following the PGD (Madry et al., 2018) algorithm to decrease the surrogate model’s cross-entropy loss that takes as inputs the adversarial examples (the poisoned samples in our case) and the target class label. This procedure is described formally in Algorithm 1. The $\bm{\delta}$ is the generated UPGD that will be used as a backdoor trigger; thus, $G_{x}({\bm{x}})={\bm{x}}+\bm{\delta}$. $S$ is the ball function with the radius $\epsilon$, and the small $\epsilon$ guarantees the imperceptibility of the backdoor trigger as it controls the perturbation’s magnitude.
@@
1:  $S=B(\bm{\delta};\epsilon)=\{\bm{\delta}\in\mathbb{R}^{d_{c}\times d_{h}\times d_{w}}:||\bm{\delta}||_{\infty}\leq\epsilon\}$
2:  $\bm{\delta}\leftarrow$ random_initialization $\land\bm{\delta}\in S$
@@
Adversarial Backdoor Injection for parameter-space stealthiness. Backdoor neurons (i.e., trigger-related neurons) regularly show higher activation values for inputs that contain the trigger, which results in powerful performance (Liu et al., 2019; Wang et al., 2022a; Lin et al., 2024). To this end, backdoor training needs to substantially increase the magnitude of parameters of backdoor neurons (Wu and Wang, 2021; Li et al., 2023a; Zheng et al., 2022), which harms the parameter-space stealthiness of backdoor attacks.
One way to find the sensitive neurons with higher activation values is to analyze the Lipschitz continuity of the network. Leveraging this fact, we introduce a novel backdoor training mechanism, Adversarial Backdoor Injection, to increase the parameter-space backdoor stealthiness. Specifically, each neuron’s Upper bound of Channel Lipschitz Condition (UCLC (Zheng et al., 2022)) is calculated, based on which the weights of these suspicious neurons are set to the mean of all neurons’ weights in the corresponding layer after every training epoch. In our implementation, we use the weights before every batch normalization as the neuron weights corresponding to the channel setting in UCLC. We prune neurons by substituting their weights with the mean ones because pruning to zeros makes the training unable to converge in our experiments. Formally, the $k_{th}$ parameter of the $l_{th}$ layer, ${\bm{\theta}}_{l}^{(k)}$, is updated as follows:
(3) ${\bm{\theta}}_{l}^{(k)}:=\left\{\begin{aligned} &\text{mean}({\bm{\theta}}_{l}),&\sigma({\bm{\theta}}_{l}^{(k)})>\text{mean}(\sigma({\bm{\theta}}_{l}))+u\times\text{std}(\sigma({\bm{\theta}}_{l}))\\ &{\bm{\theta}}_{l}^{(k)},&\text{otherwise,}\end{aligned}\right.$
@@
In adversarial training (Madry et al., 2018), adversarial examples are introduced during training to increase the model’s robustness during inference. Similarly, during the Adversarial Backdoor Injection, we use backdoor defenses to increase the resistance of backdoor attacks to parameter-space defenses. At the end of each training epoch, Adversarial Backdoor Injection prunes the trained model to decrease the weights of backdoor neurons. Iteratively, backdoored neurons spread across the whole model instead of forming a few prominent backdoor neurons, as illustrated in Figure 1.
@@
Attack No Defense FP (Liu et al., 2018a) ANP (Wu and Wang, 2021) CLP (Zheng et al., 2022) RNP (Li et al., 2023a) Average
@@
BadNets (Gu et al., 2019) 93.13 100 92.42 71.71 91.60 1.06 88.99 49.02 84.04 13.82 89.26 $\downarrow$3.87 33.90 $\downarrow$66.10
Blend (Chen et al., 2017) 94.42 100 93.08 99.99 93.57 0.33 90.3 0.54 94.63 57.98 92.89 $\downarrow$1.53 39.71 $\downarrow$60.29
WaNet (Nguyen and Tran, 2021) 93.60 99.37 92.96 4.60 91.08 0.49 91.53 2.12 92.86 3.17 92.11 $\downarrow$1.49 2.59 $\downarrow$96.78
IAD (Nguyen and Tran, 2020) 92.88 97.10 91.96 1.22 92.84 0.71 92.24 0.74 92.72 0.42 92.44 $\downarrow$0.44 0.77 $\downarrow$96.33
AdvDoor (Zhang et al., 2021) 93.97 100 93.37 98.69 91.46 28.83 89.22 6.13 90.17 44.60 91.05 $\downarrow$2.92 44.56 $\downarrow$55.44
Bpp (Wang et al., 2022b) 94.19 99.93 93.38 18.89 92.96 2.97 93.37 1.89 92.2 5.79 92.98 $\downarrow$1.21 7.39 $\downarrow$92.54
LC (Turner et al., 2019) 94.31 100 92.22 93.57 91.02 24.43 90.96 0.38 82.70 33.60 89.23 $\downarrow$5.08 37.99 $\downarrow$62.01
Narcissus (Zeng et al., 2023) 93.58 99.64 93.49 96.54 89.76 49.18 93.19 97.82 91.10 94.59 91.88 $\downarrow$1.70 84.53 $\downarrow$15.11
Adap-blend (Qi et al., 2023a) 92.74 99.67 92.06 95.50 86.48 67.73 92.49 99.62 78.63 1.56 87.42 $\downarrow$5.32 66.10 $\downarrow$33.57
SSDT (Mo et al., 2024) 93.70 90.30 93.41 0.80 93.88 0.60 93.66 1.20 93.99 3.30 93.74 $\uparrow$0.04 1.47 $\downarrow$88.83
DFST (Cheng et al., 2021) 95.23 100 94.79 93.84 94.64 3.72 92.43 3.53 93.29 12.68 93.79 $\downarrow$1.44 28.44 $\downarrow$71.56
DFBA (Cao et al., 2024) 88.99 100 86.85 0.03 88.96 9.55 88.96 9.57 88.96 0.90 88.43 $\downarrow$0.56 5.01 $\downarrow$94.99
@@
Attack vanilla FT FT-SAM (Zhu et al., 2023) I-BAU (Zeng et al., 2022) FST (Min et al., 2023) BTI-DBF(U) (Xu et al., 2024a) Average
@@
BadNets (Gu et al., 2019) 91.07 43.96 92.01 2.84 92.60 76.02 92.40 13.10 91.26 13.12 91.87 $\downarrow$1.26 29.81 $\downarrow$70.19
Blend (Chen et al., 2017) 91.64 99.61 92.52 1.73 91.84 8.84 93.40 100 91.86 100 92.25 $\downarrow$2.17 62.04 $\downarrow$37.96
WaNet (Nguyen and Tran, 2021) 91.11 0.99 90.89 1.03 87.98 0.81 92.17 0.04 90.30 4.89 90.49 $\downarrow$3.11 1.55 $\downarrow$97.82
IAD (Nguyen and Tran, 2020) 90.83 2.16 92.18 2.87 88.4 15.68 91.29 0.00 89.54 1.59 90.45 $\downarrow$2.43 4.46 $\downarrow$92.64
AdvDoor (Zhang et al., 2021) 91.25 68.68 92.18 1.23 89.29 16.99 91.06 99.99 90.25 100 90.81 $\downarrow$3.16 57.38 $\downarrow$42.62
Bpp (Wang et al., 2022b) 91.36 3.40 91.38 1.00 92.06 6.46 93.23 26.83 90.61 2.73 91.73 $\downarrow$2.46 8.08 $\downarrow$91.85
LC (Turner et al., 2019) 90.26 88.52 91.46 1.91 85.87 5.11 91.80 13.11 90.71 4.37 90.02 $\downarrow$4.29 22.60 $\downarrow$77.40
Narcissus (Zeng et al., 2023) 91.70 92.91 91.76 23.98 91.48 51.74 90.06 54.22 90.94 98.11 91.19 $\downarrow$2.39 64.19 $\downarrow$35.45
Adap-blend (Qi et al., 2023a) 92.42 98.73 91.23 22.4 85.38 37.31 90.91 1.19 89.17 7.09 89.82 $\downarrow$2.92 33.34 $\downarrow$66.33
SSDT (Mo et al., 2024) 93.74 0.70 93.15 0.60 90.27 3.10 92.85 0.20 90.79 1.40 92.16 $\downarrow$1.54 1.20 $\downarrow$89.10
DFST (Cheng et al., 2021) 95.01 2.07 94.70 0.00 89.75 19.11 93.06 2.66 90.41 22.34 92.59 $\downarrow$2.64 9.24 $\downarrow$90.76
DFBA (Cao et al., 2024) 86.68 10.10 86.03 5.24 85.48 100 82.76 57.62 84.49 100 85.09 $\downarrow$3.90 54.59 $\downarrow$45.41
@@
Datasets Attack No Defense FT-SAM (Zhu et al., 2023) I-BAU (Zeng et al., 2022) CLP (Zheng et al., 2022) Average
@@
ImageNet200 BadNets (Gu et al., 2019) 80.65 91.03 79.89 2.21 70.28 26.06 70.74 64.86 73.64 $\downarrow$7.01 31.04 $\downarrow$59.99
Blend (Chen et al., 2017) 80.70 95.63 80.19 0.39 76.13 30.81 80.02 23.38 78.78 $\downarrow$1.92 18.19 $\downarrow$77.44
WaNet (Nguyen and Tran, 2021) 81.24 99.97 80.41 0.66 75.67 47.27 77.18 99.78 77.75 $\downarrow$3.49 49.24 $\downarrow$50.73
IAD (Nguyen and Tran, 2020) 79.74 99.98 75.49 0.68 77.44 15.18 76.97 84.49 76.63 $\downarrow$3.11 33.45 $\downarrow$66.53
AdvDoor (Zhang et al., 2021) 80.72 100 79.52 98.90 74.03 61.31 77.90 100 77.15 $\downarrow$3.57 86.74 $\downarrow$13.26
Bpp (Wang et al., 2022b) 81.36 92.74 79.37 1.05 76.53 3.21 80.10 2.34 78.67 $\downarrow$2.69 2.19 $\downarrow$90.55
Narcissus (Zeng et al., 2023) 81.73 81.28 80.00 83.37 77.03 56.19 80.99 86.37 79.34 $\downarrow$2.39 75.31 $\downarrow$5.97
SSDT (Mo et al., 2024) 75.45 100 78.19 76.00 76.26 22.00 76.02 94.00 76.82 $\uparrow$1.37 64.00 $\downarrow$36.00
@@
GTSRB BadNets (Gu et al., 2019) 97.19 100 95.57 0.48 92.02 29.22 96.38 0.47 94.66 $\downarrow$2.53 10.06 $\downarrow$89.94
Blend (Chen et al., 2017) 95.92 100 93.36 0.21 92.64 38.27 93.21 0.00 93.07 $\downarrow$2.85 12.83 $\downarrow$87.17
WaNet (Nguyen and Tran, 2021) 98.69 99.77 92.18 0.45 91.25 0.00 90.14 18.14 91.19 $\downarrow$7.50 6.19 $\downarrow$93.58
IAD (Nguyen and Tran, 2020) 99.08 99.65 92.72 0.10 90.11 0.35 98.08 14.63 93.64 $\downarrow$5.44 5.03 $\downarrow$94.62
AdvDoor (Zhang et al., 2021) 95.80 99.99 93.94 32.26 92.67 38.20 90.09 66.39 92.23 $\downarrow$3.57 45.62 $\downarrow$54.37
Bpp (Wang et al., 2022b) 98.69 99.93 91.27 0.00 92.61 0.23 97.16 2.29 93.68 $\downarrow$5.01 0.84 $\downarrow$99.09
Narcissus (Zeng et al., 2023) 95.60 97.18 93.61 54.55 92.87 80.74 93.99 97.60 93.49 $\downarrow$2.11 77.63 $\downarrow$19.55
SSDT (Mo et al., 2024) 96.02 77.78 93.11 0.00 90.82 0.00 94.65 19.31 92.86 $\downarrow$3.16 6.44 $\downarrow$71.34

### v3

@@ 3.2. Lack of Parameter-Space Stealthiness
As introduced in the related work, early backdoor attacks that introduce noticeable changes in either input (Gu_2019_badnet; chen_2017_blend) or feature space (nguyen2021wanet; nguyen2020inputaware) have been empirically shown powerful, even with very low poisoning rates (Gu_2019_badnet; zeng2023narcisus). Focusing on the backdoor-introduced noticeable changes, backdoor defenses are improved to distinguish backdoor patterns in either input or feature space (wang2022featurere; lin2024unveiling). Meanwhile, backdoor attacks are optimized to increase stealthiness in input (nguyen2021wanet) or feature space (qi2022revisitingadaptive). However, regardless of the implementation of input- or feature-space attack logic, backdoor behaviors are eventually embedded in the backdoored model’s parameters. For this reason, it is important to investigate whether backdoor attacks introduce visible changes in the parameter space of the attacked models that can be used by the parameter-space defenses.
Considering this observation, we ran an initial experiment to understand the behavior of neurons in backdoored models. We use the TAC values (zheng2022clp) to quantify the relevance of a neuron to the backdoor behavior. The TAC values show the change in the output of each neuron (a feature channel of convolutional layers) before and after the trigger is attached to the input. Thus, TAC quantifies a neuron’s sensitivity to the backdoor trigger. A high TAC value indicates that the neuron is strongly related to the backdoor behavior, whereas a low TAC value shows it is not. TAC takes the exact trigger information into account when measuring the backdoor effect, which makes TAC a straightforward and effective method that captures the relevant backdoor neurons’ behaviour. Specifically, TAC is defined as:
[…]
@@ 3.3. Grond for Comprehensive Stealthiness
Backdoor trigger generation for input-space stealthiness. We use imperceptible adversarial perturbations to generate imperceptible backdoor triggers inspired by adversarial example studies (Moosavi2017uap; zhang2021advdoor). We modify the original PGD algorithm to generate a universal PGD (UPGD) perturbation as the backdoor trigger. UPGD contains non-robust but generalizable semantic information (tsipras2018robustnessoddwithacc), which correlates with the benign functions of the victim model and shortens the distance between poisoned data and the target classification region (zhang2021advdoor). Consequently, backdoor patterns tend to make fewer prominent changes to the victim network.
Similar to (zeng2023narcisus; zhang2021advdoor), UPGD is generated on a well-trained surrogate model trained on the clean training set. The architecture and parameters of the surrogate model do not necessarily need to be the same as the victim model (see Table 15 in App. C.2). UPGD is optimized following the PGD (madry2018pgd) algorithm to decrease the surrogate model’s cross-entropy loss that takes as inputs the adversarial examples (the poisoned samples in our case) and the target class label. This procedure is described formally in Algorithm 1. The $\bm{\delta}$ is the generated UPGD that will be used as a backdoor trigger; thus, $G_{x}({\bm{x}})={\bm{x}}+\bm{\delta}$. $S$ is the ball function with the radius $\epsilon$, and the small $\epsilon$ guarantees the imperceptibility of the backdoor trigger as it controls the perturbation’s magnitude.
@@
[…]
@@
Adversarial Backdoor Injection for parameter-space stealthiness. Backdoor neurons (i.e., trigger-related neurons) regularly show higher activation values for inputs that contain the trigger, which results in powerful performance (liu_2019_abs; wang2022featurere; lin2024unveiling). To this end, backdoor training needs to substantially increase the magnitude of parameters of backdoor neurons (Wu_2021_anp; li2023rnp; zheng2022clp), which harms the parameter-space stealthiness of backdoor attacks.
One way to find the sensitive neurons with higher activation values is to analyze the Lipschitz continuity of the network. Leveraging this fact, we introduce a novel backdoor training mechanism, Adversarial Backdoor Injection, to increase the parameter-space backdoor stealthiness. Specifically, each neuron’s Upper bound of Channel Lipschitz Condition (UCLC (zheng2022clp)) is calculated, based on which the weights of these suspicious neurons are set to the mean of all neurons’ weights in the corresponding layer after every training epoch. In our implementation, we use the weights before every batch normalization as the neuron weights corresponding to the channel setting in UCLC. We prune neurons by substituting their weights with the mean ones because pruning to zeros makes the training unable to converge in our experiments. Formally, the $k_{th}$ parameter of the $l_{th}$ layer, ${\bm{\theta}}_{l}^{(k)}$, is updated as follows:
(3) ${\bm{\theta}}_{l}^{(k)}=\left\{\begin{aligned} &\text{mean}({\bm{\theta}}_{l}),&\sigma({\bm{\theta}}_{l}^{(k)})>\text{mean}(\sigma({\bm{\theta}}_{l}))+u\times\text{std}(\sigma({\bm{\theta}}_{l}))\\ &{\bm{\theta}}_{l}^{(k)},&\text{otherwise,}\end{aligned}\right$.
@@
In adversarial training (madry2018pgd), adversarial examples are introduced during training to increase the model’s robustness during inference. Similarly, during the Adversarial Backdoor Injection, we use backdoor defenses to increase the resistance of backdoor attacks to parameter-space defenses. At the end of each training epoch, Adversarial Backdoor Injection prunes the trained model to decrease the weights of backdoor neurons. Iteratively, backdoored neurons spread across the whole model instead of forming a few prominent backdoor neurons, as illustrated in Figure 1.
@@
Attack No Defense FP (liu2018finepruning) ANP (Wu_2021_anp) CLP (zheng2022clp) RNP (li2023rnp) Average
@@
BadNets (Gu_2019_badnet) 93.13 100 92.42 71.71 91.60 1.06 88.99 49.02 84.04 13.82 89.26 $\downarrow$3.87 33.90 $\downarrow$66.10
Blend (chen_2017_blend) 94.42 100 93.08 99.99 93.57 0.33 90.3 0.54 94.63 57.98 92.89 $\downarrow$1.53 39.71 $\downarrow$60.29
WaNet (nguyen2021wanet) 93.60 99.37 92.96 4.60 91.08 0.49 91.53 2.12 92.86 3.17 92.11 $\downarrow$1.49 2.59 $\downarrow$96.78
IAD (nguyen2020inputaware) 92.88 97.10 91.96 1.22 92.84 0.71 92.24 0.74 92.72 0.42 92.44 $\downarrow$0.44 0.77 $\downarrow$96.33
AdvDoor (zhang2021advdoor) 93.97 100 93.37 98.69 91.46 28.83 89.22 6.13 90.17 44.60 91.05 $\downarrow$2.92 44.56 $\downarrow$55.44
Bpp (Wang_2022_bpp) 94.19 99.93 93.38 18.89 92.96 2.97 93.37 1.89 92.2 5.79 92.98 $\downarrow$1.21 7.39 $\downarrow$92.54
LC (turner2019lc) 94.31 100 92.22 93.57 91.02 24.43 90.96 0.38 82.70 33.60 89.23 $\downarrow$5.08 37.99 $\downarrow$62.01
Narcissus (zeng2023narcisus) 93.58 99.64 93.49 96.54 89.76 49.18 93.19 97.82 91.10 94.59 91.88 $\downarrow$1.70 84.53 $\downarrow$15.11
Adap-blend (qi2022revisitingadaptive) 92.74 99.67 92.06 95.50 86.48 67.73 92.49 99.62 78.63 1.56 87.42 $\downarrow$5.32 66.10 $\downarrow$33.57
SSDT (mo2024ssdt) 93.70 90.30 93.41 0.80 93.88 0.60 93.66 1.20 93.99 3.30 93.74 $\uparrow$0.04 1.47 $\downarrow$88.83
DFST (Cheng2021dfst) 95.23 100 94.79 93.84 94.64 3.72 92.43 3.53 93.29 12.68 93.79 $\downarrow$1.44 28.44 $\downarrow$71.56
DFBA (cao2024dfba) 88.99 100 86.85 0.03 88.96 9.55 88.96 9.57 88.96 0.90 88.43 $\downarrow$0.56 5.01 $\downarrow$94.99
@@
Attack vanilla FT FT-SAM (Zhu2023ftsam) I-BAU (zeng2022ibau) FST (min2023fst) BTI-DBF(U) (xu2024btidbf) Average
@@
BadNets (Gu_2019_badnet) 91.07 43.96 92.01 2.84 92.60 76.02 92.40 13.10 91.26 13.12 91.87 $\downarrow$1.26 29.81 $\downarrow$70.19
Blend (chen_2017_blend) 91.64 99.61 92.52 1.73 91.84 8.84 93.40 100 91.86 100 92.25 $\downarrow$2.17 62.04 $\downarrow$37.96
WaNet (nguyen2021wanet) 91.11 0.99 90.89 1.03 87.98 0.81 92.17 0.04 90.30 4.89 90.49 $\downarrow$3.11 1.55 $\downarrow$97.82
IAD (nguyen2020inputaware) 90.83 2.16 92.18 2.87 88.4 15.68 91.29 0.00 89.54 1.59 90.45 $\downarrow$2.43 4.46 $\downarrow$92.64
AdvDoor (zhang2021advdoor) 91.25 68.68 92.18 1.23 89.29 16.99 91.06 99.99 90.25 100 90.81 $\downarrow$3.16 57.38 $\downarrow$42.62
Bpp (Wang_2022_bpp) 91.36 3.40 91.38 1.00 92.06 6.46 93.23 26.83 90.61 2.73 91.73 $\downarrow$2.46 8.08 $\downarrow$91.85
LC (turner2019lc) 90.26 88.52 91.46 1.91 85.87 5.11 91.80 13.11 90.71 4.37 90.02 $\downarrow$4.29 22.60 $\downarrow$77.40
Narcissus (zeng2023narcisus) 91.70 92.91 91.76 23.98 91.48 51.74 90.06 54.22 90.94 98.11 91.19 $\downarrow$2.39 64.19 $\downarrow$35.45
Adap-blend (qi2022revisitingadaptive) 92.42 98.73 91.23 22.4 85.38 37.31 90.91 1.19 89.17 7.09 89.82 $\downarrow$2.92 33.34 $\downarrow$66.33
SSDT (mo2024ssdt) 93.74 0.70 93.15 0.60 90.27 3.10 92.85 0.20 90.79 1.40 92.16 $\downarrow$1.54 1.20 $\downarrow$89.10
DFST (Cheng2021dfst) 95.01 2.07 94.70 0.00 89.75 19.11 93.06 2.66 90.41 22.34 92.59 $\downarrow$2.64 9.24 $\downarrow$90.76
DFBA (cao2024dfba) 86.68 10.10 86.03 5.24 85.48 100 82.76 57.62 84.49 100 85.09 $\downarrow$3.90 54.59 $\downarrow$45.41
@@
Datasets Attack No Defense FT-SAM (Zhu2023ftsam) I-BAU (zeng2022ibau) CLP (zheng2022clp) Average
@@
ImageNet200 BadNets (Gu_2019_badnet) 80.65 91.03 79.89 2.21 70.28 26.06 70.74 64.86 73.64 $\downarrow$7.01 31.04 $\downarrow$59.99
Blend (chen_2017_blend) 80.70 95.63 80.19 0.39 76.13 30.81 80.02 23.38 78.78 $\downarrow$1.92 18.19 $\downarrow$77.44
WaNet (nguyen2021wanet) 81.24 99.97 80.41 0.66 75.67 47.27 77.18 99.78 77.75 $\downarrow$3.49 49.24 $\downarrow$50.73
IAD (nguyen2020inputaware) 79.74 99.98 75.49 0.68 77.44 15.18 76.97 84.49 76.63 $\downarrow$3.11 33.45 $\downarrow$66.53
AdvDoor (zhang2021advdoor) 80.72 100 79.52 98.90 74.03 61.31 77.90 100 77.15 $\downarrow$3.57 86.74 $\downarrow$13.26
Bpp (Wang_2022_bpp) 81.36 92.74 79.37 1.05 76.53 3.21 80.10 2.34 78.67 $\downarrow$2.69 2.19 $\downarrow$90.55
Narcissus (zeng2023narcisus) 81.73 81.28 80.00 83.37 77.03 56.19 80.99 86.37 79.34 $\downarrow$2.39 75.31 $\downarrow$5.97
SSDT (mo2024ssdt) 75.45 100 78.19 76.00 76.26 22.00 76.02 94.00 76.82 $\uparrow$1.37 64.00 $\downarrow$36.00
@@
GTSRB BadNets (Gu_2019_badnet) 97.19 100 95.57 0.48 92.02 29.22 96.38 0.47 94.66 $\downarrow$2.53 10.06 $\downarrow$89.94
Blend (chen_2017_blend) 95.92 100 93.36 0.21 92.64 38.27 93.21 0.00 93.07 $\downarrow$2.85 12.83 $\downarrow$87.17
WaNet (nguyen2021wanet) 98.69 99.77 92.18 0.45 91.25 0.00 90.14 18.14 91.19 $\downarrow$7.50 6.19 $\downarrow$93.58
IAD (nguyen2020inputaware) 99.08 99.65 92.72 0.10 90.11 0.35 98.08 14.63 93.64 $\downarrow$5.44 5.03 $\downarrow$94.62
AdvDoor (zhang2021advdoor) 95.80 99.99 93.94 32.26 92.67 38.20 90.09 66.39 92.23 $\downarrow$3.57 45.62 $\downarrow$54.37
Bpp (Wang_2022_bpp) 98.69 99.93 91.27 0.00 92.61 0.23 97.16 2.29 93.68 $\downarrow$5.01 0.84 $\downarrow$99.09
Narcissus (zeng2023narcisus) 95.60 97.18 93.61 54.55 92.87 80.74 93.99 97.60 93.49 $\downarrow$2.11 77.63 $\downarrow$19.55
SSDT (mo2024ssdt) 96.02 77.78 93.11 0.00 90.82 0.00 94.65 19.31 92.86 $\downarrow$3.16 6.44 $\downarrow$71.34